Jamal has a success rate of 68% for scoring on the free throws in basketball
a) what is the probability that he misses the basket on his 3rd throw ?
b) what is the expected waiting time before he misses the basket on free throws ?
1
Expert's answer
2020-03-13T12:18:57-0400

a) "P=0.68*0.68*(1-0.68)\approx0.1480".


b) "E(X)=\frac{1}{1-0.68}=3.125."